# Which Demographics do LLMs Default to During Annotation?

### v2

@@ 4.4 RQ4: Are observed patterns consistent across different large language models?
The patterns identified in the analysis presented in Table 3 demonstrate consistency across the two models examined, particularly in their ability to replicate human annotations. Notably, Table 4 highlighted specific behaviors of the LLMs in response to socio-demographic prompting. However, each of these distinct behaviors was only observed in one of the models. Overall, both LLMs tested in this study did not display clear default personas. Thus, the models remain consistent in this aspect.
@@ 5 Conclusion and Future Work
With this paper, we present the first analysis of the effect of socio-demographic prompting on tasks in the Popquorn data set. This corpus has been developed particularly for the study of the impact of demographic variables and therefore also offers itself for LLM-based studies.
We find that demographic prompting has some measurable effects on the annotation behaviors of large language models, contrasting with placebo prompting, which elicited no consistent changes across various attributes. Specifically, our analyses show that LLMs exhibit variations in annotation based on demographic attributes, particularly highlighting influences related to gender, race and age. While we cannot infer one concrete, unique default persona, we conclude that large language models do not represent all members of a society alike, and that socio-demographic prompting does influence the result in a structured manner. This result is in contrast to the findings by previous work that studied other corpora and other models.
Not all of our findings are overall consistent across models. While this could be considered an issue that is inherent to LLMs, it also presents itself with a substantial challenge: depending on the model that is used by an end-user, the impact of socio-demographic information varies, and different models default to different demographic information.
We therefore propose that future work should focus on the development of reliable instruments to measure the influence of both implicitly and explicitly provided demographic information about the user. We presume that making such information explicit and ensuring it is reliably interpreted by models is crucial not only to mitigate issues related to unknown biases that are not transparent to users or automatic processing pipelines, but also to create models that accurately replicate the diversity and variance across different individuals.
@@
The goal of our work is to develop an overall understanding of the impact of socio-demographic prompting on language models. Naturally, such study can only focus on a limited set of models and variables. Particularly limiting is the set of attributes available in data sets, which are also culturally biases towards the USA. Therefore it is important to conduct more such similar studies across languages, cultures, and regions. Also in our study, a considerable limitation is the unequal representation of demographic subgroups. This disparity can lead to unreliable results and may impact the conclusions drawn from our analyses, particularly regarding the influence of demographic prompting.
We also limited our experiments to only two large language models, conducting each experiment just once, for pragmatic reasons stemming from limited compute resources. A large scale analysis across a multitude of tasks would mitigate such issue.
Finally, it is important to note that the reported differences in annotation outputs are relatively small compared to the overall Likert scale range of 1 to 5. This raises questions about the practical significance of the observed effects, necessitating further exploration to understand their impact in real-world applications.
@@
Our work has the goal to make challenges transparent which the use of large language models has. A considerable limitation from an ethical perspective is that the variables we consider are not relevant across cultures in the world. The response options of the “race” variable, in particular, are specific to US American society. Other variables might be relevant in other cultures in the world that we are not aware of. We therefore suggest to conduct future studies that consider more open sets of variables that describe the users of language model-based systems.
Beyond the limitations of variable selection, we must also reflect on the implications of our findings for issues of inequality in the context of large language models. If models effectively mimic particular demographic groups, this can lead to the privileging of the viewpoints and experiences of those groups, potentially marginalizing individuals who are not adequately represented. Furthermore, if the responses are influenced by the inclusion of certain demographic information over others, it suggests that some demographic categories are represented as more “normal” or typical than others. This pattern mirrors historical inequities that persist in new technologies, revealing an unsettling continuity in the ways that cultural biases may be perpetuated.
@@ Appendix A Additional Results
In this section we provide more detailed results of our experiments.
@@ A.3 Results for Placebo Prompting
food 1468 0.25 0.22 3793 0.26 0.19
sleep 1486 0.23 0.21 3662 0.26 0.18
friends 1546 0.25 0.21 3696 0.26 0.19
@@
purple 443 0.27 0.20 1055 0.25 0.20
turquoise 367 0.23 0.20 1056 0.23 0.21
orange 405 0.23 0.21 1033 0.26 0.18
@@
brown 413 0.21 0.21 1007 0.26 0.19
Table 8: Sample sizes and mean distance scores of placebo prompting predictions in comparison to no-info prompting predictions for models GPT-4o and Claude at two rating tasks.

### v3

@@ 4.4 RQ4: Are observed patterns consistent across different large language models?
The patterns identified in the analysis presented in Table 3 demonstrate consistency across the two models examined, particularly in their ability to replicate human annotations. Notably, Table 4 highlights specific behaviors of the LLMs in response to socio-demographic prompting. However, each of these distinct behaviors is only observed in one of the models. Overall, both tested LLMs do not display clear default personas. Thus, the models remain consistent in this aspect.
@@ 5 Conclusion and Future Work
In this paper, we present an analysis of the effect of socio-demographic prompting on tasks in the Popquorn data set. Our findings show that demographic prompting exerts measurable effects on the annotation behaviors of large language models. We contrast this to placebo prompting, which elicits no consistent changes across various attributes. Specifically, our analyses reveal that LLMs show variations in annotation based on demographic attributes, particularly for gender, race, and age. While we cannot infer one concrete, unique default persona, we conclude that large language models do not represent all members of a society alike, and that socio-demographic prompting does influence the result in a structured manner. This stands in contrast to the results from previous studies, such as those by Beck et al. (2024) and Mukherjee et al. (2024), which report no consistent patterns.
Furthermore, our results echo some of the findings from Sun et al. (2025) regarding gender influences and that predictions tend to align more closely with the perceptions of White individuals. However, our analysis includes a broader range of demographic attributes. Hu and Collier (2024) observe that while persona variables produce modest improvements, their limited explanatory power aligns with our findings on demographic prompting. Similarly, Movva et al. (2024) highlight challenges in aligning LLMs with safety evaluations across demographics, underscoring the need to understand these impacts on model behavior. Together, these insights confirm the necessity of exploring demographic influences in LLM outputs.
Our study highlights biases in the annotation behavior of LLMs regarding demographics. These biases raise critical concerns about the perpetuation of racial inequities in applications of these models. The discrepancies we observe regarding different demographic groups suggest that LLMs may entrench existing biases rather than mitigate them. This reinforces societal norms that marginalize diversity, demonstrating that models struggle to accurately represent certain demographic groups.
Not all our findings are consistent across models. While this could be considered an issue that is inherent to LLMs, it also presents itself with a substantial challenge: depending on the model that is used by an end-user, the impact of socio-demographic information varies, and different models default to different demographic information.
Consequently, we advocate for future research to concentrate on developing reliable tools to measure the influence of both implicitly and explicitly provided demographic information. Making demographic information explicit and ensuring its accurate interpretation by models is crucial not only for addressing hidden biases but also for creating models that genuinely reflect the diversity and variance among individuals. Ultimately, our work underscores the necessity for bias-aware design in training models to provide equitable representation across demographic groups and to mitigate the risks of reinforcing existing societal inequities.
@@
Our study considers only a limited set of models and variables. Particularly limiting is the set of attributes available in the data set, which are also culturally biased towards the USA. Another considerable limitation is the unequal representation of demographic subgroups. We additionally only analyze a sample of the human raters to represent the text instances equally, which might lead to a skewed representation of certain demographics. However, our sample maintains a certain degree of representativeness of the demographics in Popquorn. Using a larger sample would increase robustness of our results.
While the decision to further analyze only demographics with over 100 annotated instances aimed to enhance the manageability and interpretability of our results, it may introduce concerns regarding the statistical confidence of findings, particularly for the non-binary group, which marginally meets this threshold. Fluctuations in the observed effects may also partially arise from the group’s under-representation in this evaluation.
A limitation of our analysis is the lack of post-hoc statistical testing, which could strengthen our findings. Although such tests are often omitted in regression model comparisons with few predictors, examining them could analyze whether assumptions like homoscedasticity and collinearity might have been violated. It is unlikely that the results or the conclusion of our work is influenced, however, this can be a possible factor that might have undermined the overall process.
Our experiments are conducted with only two large language models and each experiment was run just once, for pragmatic reasons stemming from limited financial resources. A large scale analysis across larger set of tasks would mitigate this issue.
Finally, the reported differences in annotation outputs are relatively small compared to the overall Likert scale range of 1–5. This raises questions about the practical significance of some of the observed effects, necessitating further exploration to understand their impact in real-world applications.
@@
Our work has the goal to make challenges transparent which the use of large language models has. A considerable limitation from an ethical perspective is that the variables we consider are not relevant across cultures in the world. The response options of the “race” variable, in particular, are specific to US American society. Other variables might be relevant in other cultures in the world that we are not aware of. We therefore suggest to conduct future studies that consider more open sets of variables that describe the diversity of users of language model-based systems.
Beyond the limitations of variable selection, our findings also have implications for issues of inequality in the context of large language models. If models effectively mimic particular demographic groups, this can lead to the privileging of the viewpoints and experiences of those groups, potentially marginalizing individuals who are not adequately represented. Furthermore, if the responses are influenced by the inclusion of certain demographic information over others, it suggests that some demographic categories are represented as more “normal” or typical than others. This pattern mirrors historical inequities that persist in new technologies, revealing an unsettling continuity in the ways that cultural biases may be perpetuated.
ChatGPT (OpenAI, 2024a) was used to gain inspiration for formulations of our initial notes for the text of some sections of this paper, as well as to find typos.
Acknowledgments
This work has been supported by the German Research Foundation (DFG) in the projects KL2869/1–2 (CEAT, project number 380093645), KL2869/5–1 (FIBISS, 438135827), KL2869/11–1 (ITEM, 513384754), KL2869/12–1 (EMCONA, 516512112), KL2869/13–1 (INPROMPT, 521755488).
@@ Appendix A Additional Results
In this section we provide additional details regarding the results of our experiments.
@@ A.3 Results for Placebo Prompting
food 1,468 0.25 0.22 3,793 0.26 0.19
sleep 1,486 0.23 0.21 3,662 0.26 0.18
friends 1,546 0.25 0.21 3,696 0.26 0.19
@@
purple 443 0.27 0.20 1,055 0.25 0.20
turquoise 367 0.23 0.20 1,056 0.23 0.21
orange 405 0.23 0.21 1,033 0.26 0.18
@@
brown 413 0.21 0.21 1,007 0.26 0.19
Table 8: Sample sizes and mean distance scores of predictions for placebo prompting (P prompts) predictions in comparison to predictions with N prompts for models GPT-4o and Claude at two rating tasks.